A point charge of +7.5μ C is located 0.35 m from a second point charge of -7.9μ C. What is the absolute magnitude of the force on each charge?
Answer #1

F = kq1q2/r^2

F = (9x10^9)(7.5x10^-6)(7.9x10^-6)/(0.35)^2 = 4.353 N
Answer #2

Force F =kq1q2 / r2

F = (9*109*7.5*10-6*7.9*10-6) / (0.35)2 = 3.96 N
Answer #3

F = kq1q2 / d^2   =9x10^9 x 7.5x 10-6 x 7.9 x 10-6 / 0.35^2 = 0.533 N